write an equation in slope-intercept form of a line conataining the point (4,-5) with a slope equal -6.
Answer by Alan3354(21608)

write an equation in slope-intercept form of a line conataining the point (4,-5) with a slope equal -6.
--------------------
Use
y-y1 = m*(x-x1) where m is the slope and (x1,y1) is the point.
y+5 = -6*(x-4)
y+5 = -6x+24
y = -6x + 19